# VersaPRM: Multi-Domain Process Reward Model via Synthetic Reasoning Data

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated significant potential in tackling complex reasoning tasks. Specifically, they can employ a step-by-step Chain-of-Thought (CoT) approach to generate more accurate and reliable solutions (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023; Madaan et al., 2023). Moreover, by using additional test-time computation, the reasoning performance of LLMs can be further enhanced (Snell et al., 2024; Yao et al., 2024).
@@ 2 Related Work
Outcome Reward and Process Reward Models.
PRMs have proven more effective than ORMs in enhancing LLM reasoning, particularly for mathematical tasks (Luo et al., 2024; Lightman et al., 2024; Sun et al., 2024). Unlike ORMs, which focus on final outcomes, PRMs provide step-by-step feedback, improving error detection in intermediate steps and multi-step tasks (Luo et al., 2024; Lightman et al., 2024; Uesato et al., 2022; Wang et al., 2024b). Techniques like OmegaPRM (Luo et al., 2024) and Math-Shepherd (Wang et al., 2024b) reduce reliance on costly human annotations, while RLHflow (Xiong et al., 2024), OpenR (Wang et al., 2024a) and ProcessBench (Zheng et al., 2024) advance PRM evaluation and training. However, the expertise of existing PRMs is mainly limited to mathematical reasoning. Our work extends the capability of PRM to multi-domain reasoning, using synthetic reasoning data for broader applicability.
@@
Test-time inference algorithms enhance LLM reasoning by adding computation during inference. AlphaCode (Li et al., 2022) shows how test-time computing boosts competitive programming performance, while Snell et al. (2024) argue that scaling inference-time compute, rather than model parameters, yields better results by adapting compute allocation to prompt difficulty. Test-time inference includes self-improvement methods like Tree of Thoughts (ToT) (Yao et al., 2024), self-verification (Weng et al., 2023), and stepwise self-evaluation (Xie et al., 2024), and external-verifier methods such as verifier reranking (Cobbe et al., 2021; Wan et al., 2024), tool feedback (Gou et al., 2024), reward-guided reasoning (Yang et al., 2024b), and multi-agent debate (Du et al., 2024). Our work uses PRMs for reranking solutions and guiding reasoning within the external verifier paradigm.
@@
Obtaining fine-grained step-wise labeling of CoTs via expert annotation is costly and time-consuming. Automated annotation methods, such as rollout-based approaches, reduce human effort but require numerous model inference, which is computationally expensive. Recent effort to mitigate these limitations include multi-rollout generation per reasoning step (Wang et al., 2024b, d) and efficiency improvements via binary search (Luo et al., 2024). Despite these advancements, the reliance on multiple model calls remains a bottleneck. Recent studies by Gilardi et al. (2023) and Fonseca & Cohen (2024) demonstrate the potential of LLMs as data labelers. In line with the work of Zheng et al. (2024), our work uses LLMs as labelers to automatically generate process reward annotations, enabling the production of synthetic data across multiple domains without high computational costs.
@@ 3.1 Score Aggregation Methods
Min-Aggregation. We use the minimum PRM step score in a CoT as the aggregated score:
@@
Last-Aggregation. We utilize the PRM score of the last step in a CoT as the aggregated score:
@@
Average-Aggregation. We employ the average PRM step score of the CoT as the aggregated score:
@@ 3.2 Inference-Time Methods
Majority Voting (MV). MV (Wang et al., 2023) is a robust baseline inference-time method that does not require a PRM. Specifically, we first sample $N$ candidate solutions to a problem from a generator. The final answer is then determined by selecting the solution that appears most frequently among these $N$ candidates:
@@
Weighted Majority Voting (WMV). This method, as used by Uesato et al. (2022), is similar to MV. We still first sample $N$ candidate solutions. However, we weight the frequencies of CoTs with identical answers by the aggregation scores. The final answer is the one with highest sum of weights:
@@
Best-of-$N$ (BoN). This method also samples $N$ candidate solutions. It then reranks them using the aggregation score from a PRM. The answer of the solution with highest score is chosen as final answer:
@@
Beam Search. This method (Snell et al., 2024) is initialized with a fixed number of beams $N$ and width $M$. The process starts by sampling $N$ initial predictions for the first reasoning step. These are ranked via the PRM’s step score, retaining the top $\frac{N}{M}$ candidates. For each retained candidate, $M$ proposals for the next step are sampled, yielding $N$ new candidates. This iterates until all beams reach solutions or a maximum iteration limit. The final prediction is selected based on the highest aggregated PRM score across steps. See Algorithm 1 for details.
Monte Carlo Tree Search (MCTS). MCTS is a search algorithm used during test-time inference (Hao et al., 2023; Wan et al., 2024) that iteratively builds a search tree to find the CoT with the highest aggregated PRM score. A detailed description is presented in Appendix B and the pseudo-code is provided in Algorithm 2.
@@ 4.1 Open-Source Math PRMs
In addition to the open-source models, two math PRMs based on open-source models are specifically trained in this work. They are denoted as LlamaPRM800K and QwenPRM800K. More details are given in Section C.1
@@ 4.3 Multi-Domain Performance of Math PRMs
We conduct comprehensive analyses on a diverse set of models. For clarity, we report results for two representative models here, with additional evaluations available in Section D.1. The first model, Math-Shepherd (Wang et al., 2024b), is trained on synthetically generated math data labeled via a rollout-based method. The second model, Qwen-2.5-Math-PRM (Zheng et al., 2024), is a best-performing open-source PRM, trained on the high-quality expert labeled PRM800K math dataset (Lightman et al., 2024).
@@ 5.1 Chain-of-Thought Generation
For the generation of CoTs, we prompt Llama-3.1-8B-Instruct to produce step-by-step reasoning for each input question. For training, we source questions from the MMLU-Pro dataset (Wang et al., 2024c), selected for its high-quality, challenging problems spanning diverse topics. From this dataset, we randomly sample up to 500 questions per domain, ensuring that it is disjoint to the subset used for evaluation. We then generate 16 CoTs for each sampled question. Post-generation, we filter out CoTs exceeding the 2,048-token limit or containing unparsable answers.
@@ 5.2 Auto-Labeling
First, we incorporate explicit step evaluation definitions, inspired by Lightman et al. (2024), into the system prompt. Steps are categorized as GOOD, OK, or BAD: BAD for incorrect, unverifiable, or irrelevant steps; GOOD for correct, verifiable, and well-aligned steps; OK for intermediate cases. Second, we also provide the reference ground-truth answer in the prompt. The full prompt is detailed in Section A.2.
@@
In total, we sample 5,750 questions from MMLU-Pro. Among the 84,098 generated CoTs that passed filtering, 36,935 were labeled as having no incorrect steps and 47,163 were labeled as having at least one (see Table 2). This dataset, denoted as MMLU-Pro-CoT-Train (Labeled), is the first open-source multi-domain reasoning dataset with step-wise labels.
To assess the quality of our auto-labeled data, we conduct a manual evaluation on a random sample of 30 questions from the dataset. For each question, we randomly select one CoT classified as entirely correct and two CoTs flagged as containing an incorrect step. We then manually validate whether the auto-labeled judgments align with our own assessments.
For the CoTs labeled as correct by the auto-labeler, we observed an agreement rate of 83% with our manual evaluations. For CoTs labeled as incorrect, the agreement rate was 70%.
Based on these results, we estimate that approximately 75% of the CoTs in the entire dataset are correctly labeled. This level of accuracy is comparable to that of manually-labeled CoT datasets, such as PRM800K (Lightman et al., 2024), which is estimated to achieve around 80% accuracy. [^1] Refer to this GitHub issue for a discussion on PRM800K’s accuracy.
@@ 5.3 Auto-Labeling Prompt Analysis
To further understand the factors influencing auto-labeling performance, we conduct an evaluation of the auto-labeling using a simplified prompt. Specifically, we remove the system prompt defining the types of reasoning steps and exclude the reference ground-truth answer from the prompt. When re-evaluating the auto-labeling quality, we observed a drastic drop in performance, with the agreement rate for CoTs labeled as correct by the original auto-labeler decreasing by over 70%, from 83% to 7%, while the agreement rate for CoTs labeled as incorrect decreased from 70% to 62%.
@@
5.4 Counterfactual Augmentation
To generate additional examples of incorrect reasoning, we explore methods for instructing an LLM to modify steps in our correct CoTs, introducing specific types of errors. We refer to this process broadly as counterfactual augmentation. However, incorporating counterfactual error steps during PRM training was not observed to significantly improve performance. Therefore, we defer specific details and experiments using counterfactual augmentation to Section A.3.
@@ 6.1 Training of Our Multi-Domain PRM
To train VersaPRM, we employ a classification head atop an LLM, optimizing with a cross-entropy loss applied to a special classification token appended at the end of each CoT step in MMLU-Pro-CoT-Train (Labeled). Detailed specifics and hyperparameters are provided in Section C.2.
We explore several training configurations, including: 1) LoRA (Hu et al., 2022) vs. full fine-tuning for efficient training, 2) a base LLM vs. a math PRM for initializing the PRM, and 3) a Qwen-based PRM vs. a Llama-based PRM for training. Comprehensive experimental results for these studies are presented in the next section. Based on those findings, our final, our final multi-domain PRM, named VersaPRM, is initialized from our LlamaPRM800K—see Appendix C for its details—fine-tuned using LoRA on our multi-domain training dataset.
@@
We first report results of the reranking methods WMV and BoN on MMLU-Pro-CoT-Eval (Unlabeled). For both methods, we adopt Min-aggregation, as it outperforms Average and Last in aggregating PRM step scores; see Section D.4 for comparison. We also include MV as a baseline.
Comparison with Math Open-Source PRMs. We evaluate our multi-domain PRM, VersaPRM, against open-source math PRMs by partitioning MMLU-Pro-CoT-Eval (Unlabeled) into three groups: 1) Math, 2) Math-adjacent, i.e., Chemistry, Computer Science, Engineering, Physics, and 3) non-Math-adjacent domains. As shown in Figure 3, our model consistently outperforms baselines in both WMV and BoN across all domain groups.
@@
Fine-tuning with synthetic multi-domain data enhances the generalizability of PRM.
For WMV, we can see the relative performance difference increase with domain distance from core mathematics. While performance of open-source math PRMs converges to the majority voting baseline in non-mathematical domains, our multi-domain PRM maintains robust generalization.
In BoN the superiority of our multi-domain PRM is even more pronounced. Unlike open-source math PRMs, which fail to surpass the baseline of MV in Math-adjacent and non-Math-adjacent domains, our model consistently surpasses it across all domain groups.
See Section D.6 for more fine-grained details where we plot WMV and BoN for every domain of MMLU-Pro-CoT-Eval (Unlabeled). The results are consistent with Figure 3, and VersaPRM outperforms math PRMs in all domains.
@@
Ablation Experiments Using Multi-Domain PRM Trained on Math Only Subset vs. Random Subset.
@@
As shown in Figure 4, VersaPRM (random subset) achieves superior performance in WMV compared to VersaPRM (Math subset). This trend holds across both Math and non-Math domains. These findings suggest two key insights. First, our PRM is not simply learning the question format but is acquiring knowledge on how to label reasoning across diverse domains. This is why training on diverse data enables better overall performance than training on same sized data in only one domain. Second, VersaPRM (random subset) also demonstrates slightly better performance in the math domain, indicating that training on a diverse dataset may facilitate positive transfer, where insights from other domains enhance reasoning in the Math domain.
@@
Experiments Using Other Training Configurations. While our final version of VersaPRM is trained from LlamaPRM800K on our synthetic data using LoRA, we also test the following training configurations on our multi-domain dataset:
@@
We evaluate the performance of math PRMs (using LlamaPRM800K) and VersaPRM with beam search and MCTS on MMLU-Pro-CoT-Eval (Unlabeled). The results over questions in all domains, presented in Figure 6, show that MCTS outperforms beam search and that they both do better than the MV baseline. Regardless of the search algorithm used, consistent with our WMN and BoN results, VersaPRM gives boosted performance over the math PRM. Details by category results are presented in Section D.7.
Figure 6: Comparison of VersaPRM and LlamaPRM800K with beam search and MCTS. Overall in the diverse domains from MMLU-Pro-CoT-Eval (Unlabeled), VersaPRM achieves better performance.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated significant potential in tackling complex reasoning tasks. Specifically, they can employ a step-by-step Chain-of-Thought (CoT) approach to generate more accurate and reliable solutions (Wei et al., 2022b; Kojima et al., 2022; Yao et al., 2023; Madaan et al., 2023). Moreover, by using additional test-time computation, the reasoning performance of LLMs can be further enhanced (Snell et al., 2024; Yao et al., 2024).
@@ 2 Related Work
Chain-of-Thought and Process Reward Models.
PRMs have proven more effective than ORMs in enhancing LLM reasoning via CoT, particularly for mathematical tasks (Luo et al., 2024; Lightman et al., 2024; Sun et al., 2024). Unlike ORMs, which focus on final outcomes, PRMs provide step-by-step feedback, improving error detection in intermediate steps and multi-step tasks (Luo et al., 2024; Lightman et al., 2024; Uesato et al., 2022; Wang et al., 2024b). Techniques like OmegaPRM (Luo et al., 2024) and Math-Shepherd (Wang et al., 2024b) reduce reliance on costly human annotations, while RLHflow (Xiong et al., 2024), OpenR (Wang et al., 2024a) and ProcessBench (Zheng et al., 2024) advance PRM evaluation and training. However, the expertise of existing PRMs is mainly limited to mathematical reasoning. This mathematical focus aligns with findings from Sprague et al. (2025), who show that CoT prompting yields its largest gains on mathematical problems, with only modest improvements in other domains. In contrast, our work demonstrates that when CoT prompting is paired with VersaPRM at test time, consistent improvements can still be achieved across several non-mathematical domains.
@@
Test-time inference algorithms enhance LLM reasoning by adding computation during inference, offering a distinct axis of scaling compared to the conventional trade-off between training larger models versus smaller, specialized architectures (Hoffmann et al., 2022). AlphaCode (Li et al., 2022) shows how test-time computing boosts competitive programming performance, while Snell et al. (2024) argue that scaling inference-time compute, rather than model parameters, yields better results by adapting compute allocation to prompt difficulty. Test-time inference includes verifier free methods—like Tree of Thought reasoning (Yao et al., 2024), self-verification (Weng et al., 2023) and stepwise self-evaluation (Xie et al., 2024)—as well as external-verifier methods such as verifier reranking (Cobbe et al., 2021; Wan et al., 2024), tool feedback (Gou et al., 2024), reward-guided reasoning (Yang et al., 2024b), and multi-agent debate (Du et al., 2024). Our work uses PRMs for reranking solutions and guiding reasoning within the external verifier paradigm.
@@
Obtaining fine-grained step-wise labeling of CoTs via expert annotation is costly and time-consuming. Automated annotation methods, such as rollout-based approaches, reduce human effort but require numerous model inference, which is computationally expensive. Recent effort to mitigate these limitations include multi-rollout generation per reasoning step (Wang et al., 2024b, d) and efficiency improvements via binary search (Luo et al., 2024). Despite these advancements, the reliance on multiple model calls remains a bottleneck. Li et al. (2024) introduce a novel arithmetical‑puzzles and show that fine‑tuning on large-scale synthetic examples yield precise multi‑step math reasoning capabilities—though its applicability outside mathematics remains unclear.
This challenge of efficiently generating supervision data mirrors the goals of knowledge distillation, where insights from a larger “teacher” model are transferred to a smaller “student” model, often by using the teacher to generate labeled training data (Hinton et al., 2015; Gu et al., 2024). Indeed, recent studies by Gilardi et al. (2023) and Fonseca & Cohen (2024) demonstrate the potential of LLMs as data labelers. The effectiveness of leveraging a larger model as a direct evaluator or labeler in such a process relies on its inherent zero-shot evaluation capabilities (Wei et al., 2022a). This use of LLM-generated feedback is conceptually similar to methods like Constitutional AI, where larger models are used to guide smaller ones (Bai et al., 2022). Inspired by these approaches and the prompting techniques of Zheng et al. (2024), our work employs LLMs as labelers for automated process reward annotation, enabling cost-effective synthetic data generation across multiple domains.
@@ 3.1 Score Aggregation Methods
Min-Aggregation.
We use the minimum PRM step score in a CoT as the aggregated score:
@@
Last-Aggregation.
We utilize the PRM score of the last step in a CoT as the aggregated score:
@@
Average-Aggregation.
We employ the average PRM step score of the CoT as the aggregated score:
@@ 3.2 Inference-Time Methods
Majority Voting (MV).
MV (Wang et al., 2023) is a robust baseline inference-time method that does not require a PRM. Specifically, we first sample $N$ candidate solutions to a problem from a generator. The final answer is then determined by selecting the solution that appears most frequently among these $N$ candidates:
@@
Weighted Majority Voting (WMV).
This method, as used by Uesato et al. (2022), is similar to MV. We sample $N$ candidate solutions. However, we weight the frequencies of CoTs with identical answers by the aggregation scores. The final answer is the one with highest sum of weights:
@@
Best-of-$N$ (BoN).
This method also samples $N$ candidate solutions. It then reranks them using the aggregation score from a PRM. The answer of the solution with highest score is chosen as final answer:
@@
Beam Search.
This method (Snell et al., 2024) is initialized with a fixed number of beams $N$ and width $M$. The process starts by sampling $N$ initial predictions for the first reasoning step. These are ranked via the PRM’s step score, retaining the top $\frac{N}{M}$ candidates. For each retained candidate, $M$ proposals for the next step are sampled, yielding $N$ new candidates. This iterates until all beams reach solutions or a maximum iteration limit. The final prediction is selected based on the highest aggregated PRM score across steps. See Algorithm 1 for details.
Monte Carlo Tree Search (MCTS).
MCTS is a search algorithm used during test-time inference (Hao et al., 2023; Wan et al., 2024) that iteratively builds a search tree to find the CoT with the highest aggregated PRM score. A detailed description is presented in Appendix C and the pseudo-code is provided in Algorithm 2.
@@ 4.1 Open-Source Math PRMs
In addition to the open-source models, two math PRMs based on open-source models are specifically trained in this work. They are denoted as LlamaPRM800K and QwenPRM800K. More details are given in Section D.1
@@ 4.3 Multi-Domain Performance of Math PRMs
We conduct comprehensive analyses on a diverse set of models. For clarity, we report results for two representative models here, with additional evaluations available in Section E.1. The first model, Math-Shepherd (Wang et al., 2024b), is trained on synthetically generated math data labeled via a rollout-based method. The second model, Qwen-2.5-Math-PRM (Zheng et al., 2024), is a best-performing open-source PRM, trained on the high-quality expert labeled PRM800K math dataset (Lightman et al., 2024).
@@ 5.1 Chain-of-Thought Generation
For the generation of CoTs, we prompt Llama-3.1-8B-Instruct to produce step-by-step reasoning for each input question. For training, we source questions from the MMLU-Pro dataset (Wang et al., 2024c), by randomly sampling up to 500 questions per domain, ensuring that it is disjoint to the subset used for evaluation. We then generate 16 CoTs for each sampled question. Post-generation, we filter out CoTs exceeding the 2,048-token limit or containing unparsable answers.
@@ 5.2 Auto-Labeling
First, we incorporate explicit step evaluation definitions, inspired by Lightman et al. (2024), into the system prompt. Steps are categorized as GOOD, OK, or BAD: BAD for incorrect, unverifiable, or irrelevant steps; GOOD for correct, verifiable, and well-aligned steps; OK for intermediate cases. Second, we also provide the ground-truth reference answer for the question whose CoT is being graded in the prompt. The full prompt is detailed in Section A.2.
@@
In total, we sample 5,750 questions from MMLU-Pro. Among the 84,098 generated CoTs that passed filtering, 36,935 were labeled as having no incorrect steps and 47,163 were labeled as having at least one (see Table 2). The CoTs generation and labeling were done using AWS Bedrock batch inference at a total cost under \$100 (USD). This dataset, denoted as MMLU-Pro-CoT-Train (Labeled), is the first open-source multi-domain reasoning dataset with step-wise labels.
To assess the quality of our auto-labeled data, we conduct a manual evaluation on a random sample of 64 questions from the dataset. For each question, we randomly select one CoT classified as entirely correct and two CoTs flagged as containing an incorrect step. We then manually validate whether the auto-labeled judgments align with our own assessments.
For the CoTs labeled as correct by the auto-labeler, we observed an agreement rate of 80% (95% CI: 0.69–0.89) with our manual evaluations. For CoTs labeled as incorrect, the agreement rate was 71% (95% CI: 0.63–0.79).
Based on these results, we estimate that approximately 75% of the CoTs in the entire dataset are correctly labeled. This level of accuracy is comparable to that of manually-labeled CoT datasets, such as PRM800K (Lightman et al., 2024), which is estimated to achieve around 80% accuracy. [^1] See the following Github issue for a discussion: https://github.com/openai/prm800k/issues/12.
@@ 5.3 Auto-Labeling Prompt Analysis
To further understand the factors influencing auto-labeling performance, we conduct an evaluation of the auto-labeling using a simplified prompt. Specifically, we remove the system prompt defining the types of reasoning steps and exclude the reference ground-truth answer from the prompt. When re-evaluating the auto-labeling quality on a random subset of 30 questions from the original sampled 64, we observe a drastic drop in performance, with the agreement rate for CoTs labeled as correct by the original auto-labeler decreasing by over 70%, to 7%, while the agreement rate for CoTs labeled as incorrect decreased from 71% to 62%. Additionally, in Section B.4, we provide an end-to-end experiment that evaluates a PRM trained on data labeled using this ablated prompt. The findings mirror our manual evaluation—the PRM trained with the ablated prompt used via BoN achieves less than half of the performance uplift over MV that the full model attains.
@@
5.4 Synthetic Data Augmentations
To further augment the dataset, we experiment with two synthetic augmentation methods. The first, which we term counterfactual augmentation, involves generating additional examples of incorrect reasoning by prompting an LLM to modify steps in correct CoTs, thereby introducing targeted errors. The second method is a straightforward rewriting approach, where an LLM rewrites a step to preserve its original meaning but alter its style.
However, incorporating augmentations from either method during PRM training did not lead to significant performance improvements. As a result, we defer the specific details and experiments on counterfactual augmentation and rewrite augmentation to Sections B.1 and B.2, respectively.
@@ 6.1 Training of Our Multi-Domain PRM
To train VersaPRM, we employ a classification head atop an LLM, optimizing with a cross-entropy loss applied to a special classification token appended at the end of each CoT step in MMLU-Pro-CoT-Train (Labeled). Detailed specifics and hyperparameters are provided in Section D.2.
We explore several training configurations, including: 1) LoRA (Hu et al., 2022) vs. full fine-tuning for efficient training, 2) a base LLM vs. a math PRM for initializing the PRM, and 3) a Qwen-based PRM vs. a Llama-based PRM for training. Comprehensive experimental results for these studies are presented in the next section. Based on those findings, our final, our final multi-domain PRM, named VersaPRM, is initialized from our LlamaPRM800K—see Appendix D for its details—fine-tuned using LoRA on our multi-domain training dataset.
@@
We first report results of the reranking methods WMV and BoN on MMLU-Pro-CoT-Eval (Unlabeled). For both methods, we adopt Min-aggregation, as it outperforms Average and Last in aggregating PRM step scores; see Section E.3 for comparison. We also include MV as a baseline.
Comparison with Math Open-Source PRMs.
We evaluate our multi-domain PRM, VersaPRM, against open-source math PRMs by partitioning MMLU-Pro-CoT-Eval (Unlabeled) into three groups: 1) Math, 2) Math-adjacent, i.e., Chemistry, Computer Science, Engineering, Physics, and 3) non-Math-adjacent domains. As shown in Figure 3, our model consistently outperforms baselines in both WMV and BoN across all domain groups.
@@
Fine-tuning with synthetic multi-domain data enhances the PRM’s performance in non-math domains.
For WMV, we can see the relative performance difference increase with domain distance from core mathematics. While performance of math PRMs converges to the majority voting baseline in non-mathematical domains, VersaPRM maintains robust performance in these other domains.
In BoN the superiority of VersaPRM is even more pronounced. Unlike math PRMs, which fail to surpass the baseline of MV in Math-adjacent and non-Math-adjacent domains, our model consistently surpasses it across all domain groups.
See Section E.5 for more fine-grained details where we plot WMV and BoN for every domain of MMLU-Pro-CoT-Eval (Unlabeled). The results are consistent with Figure 3, and VersaPRM outperforms math PRMs in all domains.
@@
Ablation Experiments Using VersaPRM Trained on Math Only Subset vs. Random Subset.
@@
As shown in Figure 4, VersaPRM (Math subset) obtains a modest lift over LlamaPRM800K—evidence that learning the in-domain question format partially helps. Far more striking is that VersaPRM (random subset) obtains markedly higher WMV accuracy across both math and non-math tasks These findings suggest two key insights. First, our PRM is not simply learning the question format but is acquiring knowledge on how to label reasoning across diverse domains. This is why training on diverse data enables better overall performance than training on same sized data in only one domain. Second, VersaPRM (random subset) also demonstrates slightly better performance in the math domain, indicating that training on a diverse dataset may facilitate positive transfer from other domains to math.
@@
Ablation Experiments using a Hold-out Domain.
In order to verify the domain-general reasoning capabilities of VersaPRM, we conduct additional ablation experiments using a hold-out domain approach. Specifically, we excluded one domain category (law or biology) from the VersaPRM training set and evaluate the model’s performance on the held-out domain. As illustrated in Figure 5, the performance of VersaPRM trained with one domain held out remains comparable to the fully trained model across both evaluated domains. Additionally, they both surpass the performance of the math PRMs. These results underscore that the generalization ability of VersaPRM is not merely due to broader coverage in the training data, but rather represents genuine domain-general reasoning capabilities. Additional results on other domains are presented in Section E.7.
Figure 5: Hold-out domain ablation experiment results. WMV performance comparison of VersaPRM trained with each of the indicated domains held out versus the fully trained VersaPRM.
Experiments Using Other Training Configurations.
While our final version of VersaPRM is trained from LlamaPRM800K on our synthetic data using LoRA, we also test the following training configurations on our multi-domain dataset:
@@
We evaluate the performance of math PRMs (using LlamaPRM800K) and VersaPRM with beam search and MCTS on MMLU-Pro-CoT-Eval (Unlabeled). The results over questions in all domains, presented in Figure 7, show that MCTS outperforms beam search and that they both do better than the MV baseline. Regardless of the search algorithm, consistent with our WMN and BoN results, VersaPRM gives boosted performance over the math PRM. It is also interesting to note that the performance of these search algorithms saturate much slower than WMN and BoN. Details by category results are presented in Section E.6.
Figure 7: Comparison of VersaPRM and LlamaPRM800K with beam search and MCTS. The x-axis compares MCTS and Beam Search in terms of computational cost for an equivalent number of generated CoT solutions. Overall in the diverse domains from MMLU-Pro-CoT-Eval (Unlabeled), VersaPRM achieves better performance.